Math 589  —  Fall 1996

Chaotic Dynamical Systems

Class: §76024  Professor: Robert Bruner
MWF 1:55  1201 FAB
116 State  577-3179 (office)

577-2479 (dept)

Office hours: MF 10:30-11:30, W 3-4, or by appt.

Text: Nonlinear Dynamics and Chaos, Steven Strogatz

Schedule: Since this is a small course, we can adapt the following schedule to the interests and abilities of the class. Thus, this is not cast in stone, but is simply intended as a guide. In particular, the dates of the exams may change.

Week  Chapters  Tests
1  1
2  2
3  3
4  4
5  Review and Test 1 Oct. 4
6  5
7  6
8  7
9  8
10  Review and Test 2 Nov. 8
11  9
12  10
13  11
14  12
Final exam: Date per uniform schedule

Time per uniform schedule

Your course grade will be based upon the average of your test grades, with the final counted as two and homework counted as one test. Please bring a blank bluebook to each exam. Each exam will be half in-class, and half takehome. Makeup exams should not be necessary in a class this size, but can be provided in the event of an unavoidable conflict, if that you contact me as early as possible, preferably before the scheduled time for the exam.

Homework will sometimes consist of longer assignments which will need to be written up somewhat like a research project. Other homework will be more like elementary math class homework, with answers which can be written in one line or one paragraph. All work you turn in, homework and take home portions of exams, should be your own work. This does not preclude discussing the material with me or with other students or professors, but explicitly disallows copying.